OpenStudy (anonymous):

create an quadratic equation with these FRACTIONS: 1/2 and 2/3

OpenStudy (anonymous):

x=1/2, x=2/3 x-1/2=0 x-2/3=0 (x-1/2)(x-2/3)=0 Foil out

OpenStudy (anonymous):

x^2-7/6x+1/3? correct?

OpenStudy (anonymous):

I didnt really solve it out but that looks about right

OpenStudy (anonymous):

okay thanks (: